# An improved formula for Wigner function and spin polarization in a decoupling relativistic fluid at local thermodynamic equilibrium

### v2

@@ Introduction.
While the first term is the LTE integral at the decoupling, the second term accounts for the dissipative terms of physical observables [12, 15]. In this work, we will confine ourselves to the calculation of observables, notably spin polarization, at LTE at the decoupling, so that we will neglect the second integral in the eq. (2).
In previous spin polarization calculations with the LTE density operator, significant geometric approximations were employed, explicitly or not [1, 2]. However, according to analytical models [37, 38, 39] and numerical simulations [40, 41, 42, 43], the decoupling hypersurface in the central region of the QGP has roughly constant proper time $\tau=\sqrt{t^{2}-z^{2}}$ and the geometric approximations in ref. [1] appear not fully appropriate. In this work we propose a new approach for evaluating quantities at LTE, which applies to hypersurfaces with arbitrary geometry; in essence, this method relies on an exchange of the order of integrations for the leading order correction of the Wigner function, which sheds light on the nature of the correction itself. Although our derivation is focused on the spin-1/2 case the method can be readily extended to particles with arbitrary spin.
@@
Plugging the eq. (Wigner function.) into the eq. (10), a compact form of the linear correction in $\Delta\beta$ to the Wigner function can be achieved after some elaboration:
@@
where $n_{\mu}={\rm d}\Sigma_{\mu}/|{\rm d}\Sigma|$ is the outward pointing unit normal vector to $\Sigma_{\rm D}$ and the sum over $\bar{y}(x,p)$ runs over all the intersections of the particle world-line with the hypersurface for a chosen space-time point $x$, including the trivial solution $y=x$. In the above equation, the operator $D_{y}$ is defined as:
@@
Note that the derivatives in (17) do not act on the operator in (18) which is evaluated over the intersections. The formula (16) includes linear terms in the gradients of $\Delta\beta$ to all orders as well as gradients of the normal vector $n$ to the hypersurface. The latter are obviously absent if the hypersurface is a hyperplane, as it is tacitly assumed in many calculations in thermal field theory.
@@
Such conditions are believed to be generally satisfied in high-energy heavy-ion collisions, yet the second one has not been carefully considered thus far.
@@
$G^{\mu\nu}(q)$ $=$ $\frac{n_{F}(x,p_{+})-n_{F}(x,p_{-})}{(2\pi)^{3}(\beta\cdot q)}\delta\left(p^{2}+\frac{q^{2}}{4}-m^{2}\right)$ (21)
@@
where the leading contribution to the denominator arises from $W^{+}_{\text{GE}}$ and the numerator depends on $\Delta W^{+}_{\rm LE}$. At the leading order the spin polarization reads:
@@
$\hskip-31.2982pt\times\sum_{\bar{y}(x,p)}\text{sgn}[p\cdot n(\bar{y})]\left[\varpi_{\rho\lambda}(\bar{y})+\frac{2n_{\rho}(\bar{y})\xi_{\lambda\alpha}(\bar{y})p^{\alpha}}{p\cdot n(\bar{y})}\right],$
@@
where $N_{p}=\int_{\Sigma_{\rm D}}\!\!{\rm d}\Sigma\cdot p\,\theta(p^{0})n_{F}(x,p)$ is the total particle number. The two terms in Eq. (23) can be identified as contributions of the thermal vorticity [29, 56, 16] $\varpi^{\mu\nu}\equiv(\partial^{\nu}\beta^{\mu}-\partial^{\mu}\beta^{\nu})/2$, and the thermal shear tensor [18, 1, 57, 2] $\xi^{\mu\nu}\equiv(\partial^{\nu}\beta^{\mu}+\partial^{\mu}\beta^{\nu})/2$, respectively, while the curvature of $\Sigma_{\rm D}$ does not contribute at this order.
@@
The equation (23) improves upon the formulae known in the literature in a threefold manner. First, it gives a new expression for the shear-induced polarization. In contrast to previous derivations, the factor $\hat{t}^{\mu}/p\cdot\hat{t}$ in refs. [18, 1] or $u^{\mu}/p\cdot u$ in refs. [57, 2] (where $\hat{t}^{\mu}$ is the unit time vector in the QGP frame and $u^{\mu}$ is the fluid velocity), obtained as the result of significant geometric approximations, is replaced by $n^{\mu}/|p\cdot n|$. Remarkably, in the eq. (23), the vorticity and the shear induced terms both receive contributions from the derivative of the $\beta$ field along $n_{\mu}$, but in the combination these contributions cancel out. Hence, when the decoupling hypersurface is isothermal and the direction of the normal vector $n_{\mu}$ coincides with $\partial_{\mu}T$, derivatives of the temperature do not contribute. Therefore, in this upgraded formalism, the absence of temperature gradients at all orders of the expansion [18, 58, 59] arises naturally. Secondly, the thermal vorticity contribution acquires a mofidication by the sign function $\text{sgn}(p\cdot n)$, which is apparently necessary when $\Sigma_{\rm D}$ is not entirely space-like and future oriented.
@@
Finally, the formula (23) features additional contributions from the non-trivial solution of eq. (15) with $\bar{y}(x,p)\neq x$. Apparently, they can be interpreted as the residual effect of a quantum interference between particles emitted from different points of the hypersurface, the interference yielding a non-vanishing result only for the classical path from $\bar{y}(x,p)$ to $x$. However, these additional contributions are characterized by a different sign of ${\rm d}\Sigma\cdot p$ and $p\cdot n(\bar{y})$ in the formula (23) and are geometrically related to particles moving inward across the hypersurface and traversing the fluid region [60, 61, 62, 63]. Classically, this appears un-physical and it may be, in our formalism, a spurious effect generated by not having used interacting fields in the Wigner operator when dealing with in-plasma region in the formula (5). It is possible that such additional terms are thus to be discarded and a usual way to do this for the classical Cooper-Frye formula is to introduce a cutoff $\theta(p\cdot n)$ [61, 62, 63, 64]. Likewise, introducing $\theta(p\cdot n(x))\theta(p\cdot n(\bar{y}))$ in eq. (23) will remove the additional intersections in eq. (23).
@@ Acknowledgments.
The authors thank G.-L. Ma and Q. Wang for insightful discussions. This work is supported in part by the Italian Ministry of University and Research, project PRIN2022 “Advanced probes of the Quark Gluon Plasma”, Next Generation EU, Mission 4 Component 1.
@@
$\Psi(x)=\frac{1}{\sqrt{(2\pi)^{3}}}\sum_{s}\int\frac{{\rm d}^{3}{\bf{\rm p}}}{2\varepsilon}\left(u_{s}(p)\widehat{a}_{s}(p){\rm e}^{-ip\cdot x}+v_{s}(p)\widehat{b}_{s}v_{s}(p){\rm e}^{ip\cdot x}\right)\,$, (24)
@@
and likewise for antiparticles. Replacing the integral measure in eq. (24) with:
@@
$\widehat{W}^{+}(x,p)$ $=$ $\frac{1}{(2\pi)^{3}}\sum_{r,s}\int{\rm d}^{4}p_{1}\,{\rm d}^{4}p_{2}\;\delta(p_{1}^{2}-m^{2})\theta(p_{1}^{0})\delta(p_{2}^{2}-m^{2})\theta(p_{2}^{0})\,\delta\left(p-\frac{p_{1}+p_{2}}{2}\right)$ (25)
@@
$P\equiv\frac{p_{1}+p_{2}}{2}\,,\qquad q\equiv p_{1}-p_{2}\qquad\implies\qquad p_{1}=P+q/2\,,\qquad p_{2}=P-q/2$ (26)
@@
Integrating in ${\rm d}^{4}P$ is now straightforward because of the $\delta^{4}(p-P)$ in the equation (25) and, after replacing all variables according to the (26) one readily gets the (6).
@@
$\Delta W_{\text{LE}}^{+}(x,p)$ $=$ $-\frac{1}{2(2\pi)^{3}}\sum_{r,s}\int{\rm d}^{4}q\;\theta(p_{+}^{0})\theta(p_{-}^{0})e^{iq\cdot x}\delta(p\cdot q)\int_{\Sigma_{\rm D}}\!\!{\rm d}\Sigma_{\mu}(y)\int_{0}^{1}{\rm d}z\;\Delta\beta_{\nu}(y,x)\,\delta\left(p^{2}+\frac{q^{2}}{4}-m^{2}\right)$ (27)
@@
$\left\langle\widehat{a}^{\dagger}(p_{+})\widehat{a}(p_{-}),{\rm e}^{-z\beta(x)\cdot\widehat{P}}{\widehat{T}}^{\mu\nu}(y){\rm e}^{z\beta(x)\cdot\widehat{P}}\right\rangle_{c,{\rm GE}}=\langle{\rm e}^{-C\cdot{\widehat{P}}}\widehat{a}^{\dagger}(p_{+})\widehat{a}(p_{-}){\rm e}^{C\cdot{\widehat{P}}},{\widehat{T}}^{\mu\nu}(0)\rangle_{c,{\rm GE}},$ (28)
@@
thus the (28) becomes:
$\left\langle\widehat{a}^{\dagger}(p_{+})\widehat{a}(p_{-}),{\rm e}^{-z\beta(x)\cdot\widehat{P}}{\widehat{T}}^{\mu\nu}(y){\rm e}^{z\beta(x)\cdot\widehat{P}}\right\rangle_{c,{\rm GE}}={\rm e}^{-iy\cdot q}{\rm e}^{z\beta(x)\cdot q}\langle\widehat{a}^{\dagger}(p_{+})\widehat{a}(p_{-}),{\widehat{T}}^{\mu\nu}(0)\rangle_{c,{\rm GE}}\,.$ (29)
Plugging the (29) and completing the integral over $z$ leads to the eq. (Wigner function.) in the main text.
@@
$G^{\mu\nu}(q)=\sum_{N=0}^{\infty}\frac{1}{N!}\left.\left[\partial^{q}_{\nu_{1}}\partial^{q}_{\nu_{2}}\cdots\partial^{q}_{\alpha_{N}}G^{\mu\nu}(q)\right]\right|_{q^{\mu}=0}q^{\nu_{1}}q^{\nu_{2}}\cdots q^{\nu_{N}}\,.$ (30)
@@
$\Delta W_{\text{LE}}^{+}(x,p)$ $=$ $\frac{1}{|p^{0}|}\sum_{N=0}^{\infty}\frac{(-i)^{N}}{N!}\left.\left[\partial^{q}_{\nu_{1}}\partial^{q}_{\nu_{2}}\cdots\partial^{q}_{\nu_{N}}G^{\mu\nu}(q)\right]\right|_{q=0}$ (31)
@@
acting only on the $\delta$-function and $\Delta\beta_{\nu}(y,x)$, respectively. Thereby, the partial derivatives in eq. (31) can be recast as $\left[\partial^{q}\cdot\partial^{(1)}_{x}\right]^{N}$ and taken out of the integral. The latter can be expressed as follows, by means of the Binomial theorem:
@@
$\delta^{3}\left({\bf y}-{\bf x}-\frac{\bf p}{p^{0}}(y^{0}-x^{0})\right)=\sum_{i}\frac{|p^{0}|}{|p\cdot\sigma(\bar{y}_{k,i})|}\delta^{3}({\bf y}-\bar{{\bf y}}_{k,i})\,$, (32)
@@
$\Delta W_{\text{LE}}^{+}(x,p)$ $=$ $\sum_{N=0}^{\infty}\sum_{k,i}s_{k}\frac{(-1)^{N}}{N!}\left.\left[\partial^{q}_{\nu_{1}}\ldots\partial^{q}_{\nu_{N}}G^{\mu\nu}(q)\right]\right|_{q=0}$ (33)
@@
where $s_{k}$ is a sign, which is +1 if $\sigma_{\mu}$ has the same direction of the outward pointing vector ${\rm d}\Sigma_{\mu}$ and 1-1 otherwise. In eq. (33) we have introduced the total derivative:
@@
$\frac{{\rm d}}{{\rm d}x^{\mu}}g(x,\bar{y}_{k,i}(x))=\frac{\partial}{\partial x^{\mu}}g(x,\bar{y}_{k,i}(x))+\frac{\partial\bar{y}_{k,i}^{\nu}(x)}{\partial x^{\mu}}\frac{\partial g(x,y)}{\partial y^{\nu}}\Bigg|_{y=\bar{y}_{k,i}(x)}$ (34)
@@
$\frac{\partial\bar{y}_{k,i}^{j}}{\partial x^{\mu}}-\frac{p^{j}}{p^{0}}\frac{\partial\bar{y}_{k,i}^{0}}{\partial\bar{y}_{k,i}^{l}}\frac{\partial\bar{y}_{k,i}^{l}}{\partial x^{\mu}}=\left[\delta^{j}_{l}+\frac{p^{j}}{p^{0}}\sigma_{l}\right]\frac{\partial\bar{y}_{k,i}^{l}}{\partial x^{\mu}}=\delta_{\mu}^{j}-\frac{p^{j}}{p^{0}}\delta_{\mu}^{0}$ (36)
@@
$\sigma_{\mu}(\bar{y}_{k,i})=\left(1,-\frac{\partial f_{k}(\bar{{\bf y}}_{k,i})}{\partial\bar{{\bf y}}_{k,i}}\right)\,.$ (37)
@@
$\frac{\partial\bar{y}_{k,i}^{j}}{\partial x^{\mu}}=\delta^{j}_{\mu}-\frac{p^{j}\sigma_{\mu}}{p\cdot\sigma}\,.$ (38)
@@
$\frac{\partial\bar{y}_{k,i}^{0}}{\partial x^{0}}=\sigma_{0}-\frac{p^{0}}{p\cdot\sigma}\qquad\qquad\frac{\partial\bar{y}_{k,i}^{0}}{\partial x^{m}}=-\frac{p^{0}\sigma_{m}}{p\cdot\sigma}$ (39)
The equations (38) and (39) can be written in a compact form as:
$\frac{\partial\bar{y}_{k,i}^{\nu}}{\partial x^{\mu}}=\delta_{\mu}^{\nu}-\frac{p^{\nu}\sigma_{\mu}(\bar{y}_{k,i})}{p\cdot\sigma(\bar{y}_{k,i})}\equiv\Delta_{\mu}^{\;\nu}(\bar{y}_{k,i})$ (40)
so that eq. (34) can be rewritten as:
$\frac{{\rm d}}{{\rm d}x^{\mu}}g(x,\bar{y}_{k,i}(x))=\frac{\partial}{\partial x^{\mu}}g(x,\bar{y}_{k,i}(x))+\Delta^{\ \nu}_{\mu}(\bar{y}_{k,i})\frac{\partial g(x,y)}{\partial y^{\nu}}\Bigg|_{y=\bar{y}_{k,i}(x)}$ (41)
By using the last result, the eq. (33) can be rewritten as:
$\Delta W_{\text{LE}}^{+}(x,p)$ $=$ $\sum_{N=0}^{\infty}\sum_{k,i}s_{k}\frac{(-1)^{N}}{N!}\left.\left[\partial^{q}_{\nu_{1}}\cdots\partial^{q}_{\nu_{N}}G^{\mu\nu}(q)\right]\right|_{q=0}\sum_{M=0}^{N}\frac{N!(-1)^{M}}{M!(N-M)!}\left[\partial^{\nu_{M+1}}_{x}+\Delta^{\alpha_{M+1}\rho_{M+1}}(\bar{y}_{k,i})\partial_{\rho_{M+1}}^{y}\right]$ (42)
@@
$\Delta W_{\text{LE}}^{+}(x,p)=\sum_{N=0}^{\infty}\sum_{k,i}s_{k}\frac{(-i)^{N}}{N!}\left.\left[\Delta^{\alpha\rho}(\bar{y}_{k,i})\partial_{\rho}^{y}\partial_{\alpha}^{q}\right]^{N}G^{\mu\nu}(p,q)\frac{\sigma_{\mu}(y)}{|p\cdot\sigma(y)|}\Delta\beta_{\nu}(y,x)\right|_{q=0,y=\bar{y}_{k,i}}\;.$ (43)
@@
$\text{tr}\left[W_{\rm GE}^{+}(x,p)\right]=\frac{4m}{(2\pi)^{3}}\delta(p^{2}-m^{2})\theta(p^{0})n_{F}(x,p)\,.$ (45)

### v3

@@ Introduction.
While the first term is the LTE integral at the decoupling, the second term accounts for the dissipative terms of physical observables [65, 10]. In this work, we will confine ourselves to the calculation of observables, notably spin polarization, at LTE at the decoupling. The dissipative terms are assumed to be small compared to the LTE part, so that we will neglect the second integral in the eq. (2).
In previous spin polarization calculations with the LTE density operator, significant geometric approximations were employed, explicitly or not [12, 41]; for instance, in ref. [12] for the purpose of calculating the LTE contribution, the region encompassed by the initial and the decoupling hypersurface was approximated with a rectangular region with infinitely distant time-like boundaries, while in ref. [41] the LTE hypersurface should be perpendicular to the four-velocity field, which is not the case in general. In fact, according to analytical models [19, 43, 50] and numerical simulations [57, 52, 46, 28], the decoupling hypersurface does not meet either case. In this work we propose a new approach for evaluating quantities at LTE, which applies to hypersurfaces with arbitrary geometry; in essence, this method relies on an exchange of the order of integrations for the leading order correction of the Wigner function, which sheds light on the nature of the correction itself. Although our derivation is focused on the spin-1/2 case the method can be readily extended to particles with arbitrary spin.
@@
Plugging the eq. (Wigner function.) into the eq. (Wigner function.), a compact form of the linear correction in $\Delta\beta$ to the Wigner function can be achieved after some elaboration:
@@
where $n_{\mu}=s_{k}\sigma_{\mu}/\sqrt{|\sigma\cdot\sigma|}$ is the outward pointing unit normal vector to $\Sigma_{\rm D}$ and the sum over $\bar{y}(x,p)$ runs over all the intersections of the particle world-line with the hypersurface for a chosen space-time point $x$, including the trivial solution $y=x$. In the above equation, the operator $D_{y}$ is defined as:
@@
Note that the derivatives in (17) do not act on the operator in (18) which is evaluated over the intersections. The formula (16) is the linear response to $\Delta\beta$, which contains all orders of gradients of $\beta_{\nu}$ as well as gradients of the normal vector $n$ to the hypersurface. The latter are obviously absent if the hypersurface is a hyperplane, as it is tacitly assumed in many calculations in thermal field theory.
@@
Such conditions are believed to be generally satisfied in high-energy heavy-ion collisions, yet the second one has not been carefully considered thus far. Here we provide a quantitative estimates for a hypersurface with constant proper time $\tau$. The normal vector is approximated as $n^{\mu}=(\cosh\xi,\,0,\,0,\,\sinh\xi)$, with $\xi$ being the spacetime rapidity. The typical scale of its curvature is $L_{G}\sim\tau\cosh^{-1}\xi$. For most central Au+Au collisions at $\sqrt{s_{\rm NN}}=200$ GeV, $\tau\sim 10$ fm/c [51], corresponding to $6.5\,\text{fm}\leq L_{G}\leq 10\,\text{fm}$ within the interval $|\xi|<1$, which is much larger than the typical microscopic scale $l_{c}\sim 1/T\sim 1.2$ fm when $T=160$ MeV. In the peripheral region with a larger $\xi$, the hypersurface becomes more curved with a larger $l_{c}/L_{G}$, but its contribution to observables of particles at mid-rapidity is strongly suppressed by the Fermi-Dirac distribution.
@@
$G^{\mu\nu}(q)$ $=$ $\frac{n_{F}(x,p_{+})-n_{F}(x,p_{-})}{4(2\pi)^{3}(\beta\cdot q)}\delta\left(p^{2}+\frac{q^{2}}{4}-m^{2}\right)$ (21)
@@
where the leading contribution to the denominator arises from $W^{+}_{\text{GE}}$ and the numerator depends on $\Delta W^{+}_{\rm LE}$. For the convenience of studies at low-energy collisions, we recover the contribution of finite chemical potential by taking replacements $\widehat{T}^{\mu\nu}\rightarrow\widehat{j}^{\mu}$ and $\Delta\beta_{\nu}\rightarrow\Delta\zeta(y,x)\equiv\zeta(y)-\zeta(x)$ in eq. (Wigner function.). The hypersurface integral is then completed in the same way as discussed above. At the leading order the spin polarization reads:
@@
$\times\sum_{\bar{y}(x,p)}\text{sgn}[p\cdot n(\bar{y})]\bigg\{\varpi_{\rho\lambda}(\bar{y})$
@@
$+\frac{2n_{\rho}(\bar{y})}{p\cdot n(\bar{y})}\left[\xi_{\lambda\alpha}(\bar{y})p^{\alpha}{-\partial_{\lambda}^{\bar{y}}\zeta(\bar{y})}\right]\bigg\},$
@@
where $N_{p}=\int_{\Sigma_{\rm D}}\!\!{\rm d}\Sigma\cdot p\,\theta(p^{0})n_{F}(x,p)$ is the total particle number. The three terms in eq. (23) can be identified as contributions of the thermal vorticity [37, 14, 13] $\varpi^{\mu\nu}\equiv(\partial^{\nu}\beta^{\mu}-\partial^{\mu}\beta^{\nu})/2$, the thermal shear tensor [11, 12, 30, 41] $\xi^{\mu\nu}\equiv(\partial^{\nu}\beta^{\mu}+\partial^{\mu}\beta^{\nu})/2$, and the spin Hall effect [40], respectively, while the curvature of $\Sigma_{\rm D}$ does not contribute at this order.
@@
The equation (23) improves upon the formulae known in the literature in a twofold manner. First, it gives a new expression for the shear-induced polarization. In contrast to previous derivations, the factor $\hat{t}^{\mu}/p\cdot\hat{t}$ in refs. [11, 12] or $u^{\mu}/p\cdot u$ in refs. [30, 41] (where $\hat{t}^{\mu}$ is the unit time vector in the QGP frame and $u^{\mu}$ is the fluid velocity), obtained as the result of significant geometric approximations, is replaced by $n^{\mu}/|p\cdot n|$. The new formula comes down to ref. [1] if $\Sigma_{\rm D}$ is a hyperplane, and to ref. [2] if the velocity field is normal to $\Sigma_{\rm D}$, hence it reproduces previously obtained theoretical results under specific geometric assumptions. This result solves the long-standing issue understand of the ambiguity of shear-induced polarization formula, which was addressed in several papers [42, 6]. Remarkably, in the eq. (23), the vorticity and the shear induced terms both receive contributions from the derivative of the $\beta$ field along $n_{\mu}$, but in the combination these contributions cancel out. Hence, when the decoupling hypersurface is isothermal and the direction of the normal vector $n_{\mu}$ coincides with $\partial_{\mu}T$, derivatives of the temperature do not contribute. Therefore, in this upgraded formalism, the absence of temperature gradients at all orders of the expansion [11, 9, 8] arises naturally. Secondly, the thermal vorticity contribution acquires a mofidication by the sign function $\text{sgn}(p\cdot n)$, which is apparently necessary when $\Sigma_{\rm D}$ is not entirely space-like and future oriented.
@@
The formula (23) features additional contributions from the non-trivial solution of eq. (15) with $\bar{y}(x,p)\neq x$. They are geometrically related to particles moving inward across the hypersurface and traversing the fluid region [44, 20, 7, 31]. This appears un-physical and it may be, in our formalism, a spurious effect generated by not having used interacting fields in the Wigner operator when dealing with in-plasma region in the formula (5). Such additional terms need careful studies in the future but here we propose that they can be discarded as follows: for the classical Cooper-Frye formula, the usual way to drop the inward-moving particles is to introduce a cutoff $\theta(p\cdot n)$ [20, 7, 31, 33]. Likewise, introducing $\theta(p\cdot n(x))\theta(p\cdot n(\bar{y}))$ in eq. (23) will remove the additional intersections in eq. (23).
@@ Acknowledgments.
The authors thank G.-L. Ma and Q. Wang for insightful discussions. This work is supported in part by the Italian Ministry of University and Research, project PRIN2022 “Advanced probes of the Quark Gluon Plasma”, Next Generation EU, Mission 4 Component 1. The data that support the findings of this article are openly available [54].
@@
$\Psi(x)=\frac{1}{\sqrt{(2\pi)^{3}}}\sum_{s}\int\frac{{\rm d}^{3}{\bf{\rm p}}}{2\varepsilon}\left(u_{s}(p)\widehat{a}_{s}(p){\rm e}^{-ip\cdot x}+v_{s}(p)\widehat{b}_{s}v_{s}(p){\rm e}^{ip\cdot x}\right)\,$, (S1)
@@
and likewise for antiparticles. Replacing the integral measure in eq. (S1) with:
@@
$\widehat{W}^{+}(x,p)$ $=$ $\frac{1}{(2\pi)^{3}}\sum_{r,s}\int{\rm d}^{4}p_{1}\,{\rm d}^{4}p_{2}\;\delta(p_{1}^{2}-m^{2})\theta(p_{1}^{0})\delta(p_{2}^{2}-m^{2})\theta(p_{2}^{0})\,\delta\left(p-\frac{p_{1}+p_{2}}{2}\right)$ (S2)
@@
$P\equiv\frac{p_{1}+p_{2}}{2}\,,\qquad q\equiv p_{1}-p_{2}\qquad\implies\qquad p_{1}=P+q/2\,,\qquad p_{2}=P-q/2$ (S3)
@@
Integrating in ${\rm d}^{4}P$ is now straightforward because of the $\delta^{4}(p-P)$ in the equation (S2) and, after replacing all variables according to the (S3) one readily gets the (6).
@@
$\Delta W_{\text{LE}}^{+}(x,p)$ $=$ $-\frac{1}{2(2\pi)^{3}}\sum_{r,s}\int{\rm d}^{4}q\;\theta(p_{+}^{0})\theta(p_{-}^{0})e^{iq\cdot x}\delta(p\cdot q)\int_{\Sigma_{\rm D}}\!\!{\rm d}\Sigma_{\mu}(y)\int_{0}^{1}{\rm d}z\;\Delta\beta_{\nu}(y,x)\,\delta\left(p^{2}+\frac{q^{2}}{4}-m^{2}\right)$ (S4)
@@
$\left\langle\widehat{a}^{\dagger}(p_{+})\widehat{a}(p_{-}),{\rm e}^{-z\beta(x)\cdot\widehat{P}}{\widehat{T}}^{\mu\nu}(y){\rm e}^{z\beta(x)\cdot\widehat{P}}\right\rangle_{c,{\rm GE}}=\langle{\rm e}^{-C\cdot{\widehat{P}}}\widehat{a}^{\dagger}(p_{+})\widehat{a}(p_{-}){\rm e}^{C\cdot{\widehat{P}}},{\widehat{T}}^{\mu\nu}(0)\rangle_{c,{\rm GE}},$ (S5)
@@
thus the (S5) becomes:
$\left\langle\widehat{a}^{\dagger}(p_{+})\widehat{a}(p_{-}),{\rm e}^{-z\beta(x)\cdot\widehat{P}}{\widehat{T}}^{\mu\nu}(y){\rm e}^{z\beta(x)\cdot\widehat{P}}\right\rangle_{c,{\rm GE}}={\rm e}^{-iy\cdot q}{\rm e}^{z\beta(x)\cdot q}\langle\widehat{a}^{\dagger}(p_{+})\widehat{a}(p_{-}),{\widehat{T}}^{\mu\nu}(0)\rangle_{c,{\rm GE}}\,.$ (S6)
Plugging the (S6) and completing the integral over $z$ leads to the eq. (Wigner function.) in the main text.
@@
$G^{\mu\nu}(q)=\sum_{N=0}^{\infty}\frac{1}{N!}\left.\left[\partial^{q}_{\nu_{1}}\partial^{q}_{\nu_{2}}\cdots\partial^{q}_{\alpha_{N}}G^{\mu\nu}(q)\right]\right|_{q^{\mu}=0}q^{\nu_{1}}q^{\nu_{2}}\cdots q^{\nu_{N}}\,.$ (S7)
@@
$\Delta W_{\text{LE}}^{+}(x,p)$ $=$ $\frac{1}{|p^{0}|}\sum_{N=0}^{\infty}\frac{(-i)^{N}}{N!}\left.\left[\partial^{q}_{\nu_{1}}\partial^{q}_{\nu_{2}}\cdots\partial^{q}_{\nu_{N}}G^{\mu\nu}(q)\right]\right|_{q=0}$ (S8)
@@
acting only on the $\delta$-function and $\Delta\beta_{\nu}(y,x)$, respectively. Thereby, the partial derivatives in eq. (S8) can be recast as $\left[\partial^{q}\cdot\partial^{(1)}_{x}\right]^{N}$ and taken out of the integral. The latter can be expressed as follows, by means of the Binomial theorem:
@@
$\delta^{3}\left({\bf y}-{\bf x}-\frac{\bf p}{p^{0}}(y^{0}-x^{0})\right)=\sum_{i}\frac{|p^{0}|}{|p\cdot\sigma(\bar{y}_{k,i})|}\delta^{3}({\bf y}-\bar{{\bf y}}_{k,i})\,$, (S9)
@@
$\Delta W_{\text{LE}}^{+}(x,p)$ $=$ $\sum_{N=0}^{\infty}\sum_{k,i}s_{k}\frac{(-1)^{N}}{N!}\left.\left[\partial^{q}_{\nu_{1}}\ldots\partial^{q}_{\nu_{N}}G^{\mu\nu}(q)\right]\right|_{q=0}$ (S10)
@@
where $s_{k}$ is a sign, which is +1 if $\sigma_{\mu}$ has the same direction of the outward pointing vector ${\rm d}\Sigma_{\mu}$ and 1-1 otherwise. In eq. (S10) we have introduced the total derivative:
@@
$\frac{{\rm d}}{{\rm d}x^{\mu}}g(x,\bar{y}_{k,i}(x))=\frac{\partial}{\partial x^{\mu}}g(x,\bar{y}_{k,i}(x))+\frac{\partial\bar{y}_{k,i}^{\nu}(x)}{\partial x^{\mu}}\frac{\partial g(x,y)}{\partial y^{\nu}}\Bigg|_{y=\bar{y}_{k,i}(x)}$ (S11)
@@
$\frac{\partial\bar{y}_{k,i}^{j}}{\partial x^{\mu}}-\frac{p^{j}}{p^{0}}\frac{\partial\bar{y}_{k,i}^{0}}{\partial\bar{y}_{k,i}^{l}}\frac{\partial\bar{y}_{k,i}^{l}}{\partial x^{\mu}}=\left[\delta^{j}_{l}+\frac{p^{j}}{p^{0}}\sigma_{l}\right]\frac{\partial\bar{y}_{k,i}^{l}}{\partial x^{\mu}}=\delta_{\mu}^{j}-\frac{p^{j}}{p^{0}}\delta_{\mu}^{0}$ (S13)
@@
$\sigma_{\mu}(\bar{y}_{k,i})=\left(1,-\frac{\partial f_{k}(\bar{{\bf y}}_{k,i})}{\partial\bar{{\bf y}}_{k,i}}\right)\,.$ (S14)
@@
$\frac{\partial\bar{y}_{k,i}^{j}}{\partial x^{\mu}}=\delta^{j}_{\mu}-\frac{p^{j}\sigma_{\mu}}{p\cdot\sigma}\,.$ (S15)
@@
$\frac{\partial\bar{y}_{k,i}^{0}}{\partial x^{0}}=\sigma_{0}-\frac{p^{0}}{p\cdot\sigma}\qquad\qquad\frac{\partial\bar{y}_{k,i}^{0}}{\partial x^{m}}=-\frac{p^{0}\sigma_{m}}{p\cdot\sigma}$ (S16)
The equations (S15) and (S16) can be written in a compact form as:
$\frac{\partial\bar{y}_{k,i}^{\nu}}{\partial x^{\mu}}=\delta_{\mu}^{\nu}-\frac{p^{\nu}\sigma_{\mu}(\bar{y}_{k,i})}{p\cdot\sigma(\bar{y}_{k,i})}\equiv\Delta_{\mu}^{\;\nu}(\bar{y}_{k,i})$ (S17)
so that eq. (S11) can be rewritten as:
$\frac{{\rm d}}{{\rm d}x^{\mu}}g(x,\bar{y}_{k,i}(x))=\frac{\partial}{\partial x^{\mu}}g(x,\bar{y}_{k,i}(x))+\Delta^{\ \nu}_{\mu}(\bar{y}_{k,i})\frac{\partial g(x,y)}{\partial y^{\nu}}\Bigg|_{y=\bar{y}_{k,i}(x)}$ (S18)
By using the last result, the eq. (S10) can be rewritten as:
$\Delta W_{\text{LE}}^{+}(x,p)$ $=$ $\sum_{N=0}^{\infty}\sum_{k,i}s_{k}\frac{(-1)^{N}}{N!}\left.\left[\partial^{q}_{\nu_{1}}\cdots\partial^{q}_{\nu_{N}}G^{\mu\nu}(q)\right]\right|_{q=0}\sum_{M=0}^{N}\frac{N!(-1)^{M}}{M!(N-M)!}\left[\partial^{\nu_{M+1}}_{x}+\Delta^{\alpha_{M+1}\rho_{M+1}}(\bar{y}_{k,i})\partial_{\rho_{M+1}}^{y}\right]$ (S19)
@@
$\Delta W_{\text{LE}}^{+}(x,p)=\sum_{N=0}^{\infty}\sum_{k,i}s_{k}\frac{(-i)^{N}}{N!}\left.\left[\Delta^{\alpha\rho}(\bar{y}_{k,i})\partial_{\rho}^{y}\partial_{\alpha}^{q}\right]^{N}G^{\mu\nu}(p,q)\frac{\sigma_{\mu}(y)}{|p\cdot\sigma(y)|}\Delta\beta_{\nu}(y,x)\right|_{q=0,y=\bar{y}_{k,i}}\;.$ (S20)
@@
$\text{tr}\left[W_{\rm GE}^{+}(x,p)\right]=\frac{4m}{(2\pi)^{3}}\delta(p^{2}-m^{2})\theta(p^{0})n_{F}(x,p)\,.$ (S22)